Circular No. 6024

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) The Hubble Space Telescope Team reports a location for the feature

created by the impact of fragment A = 21. A streak ranges from jovicentric latitude B = −42^{o} ± 1^{o}, System III longitude L = 186^{o} to B = −39^{o} ± 1^{o}, L = 188^{o}. At 410 nm the maximum radial extent of the dark region surrounding the streak is 12 000 km. Although the region appears to be spatially circularly symmetric, the intensity distribution appears to be asymmetric, being more apparent to the south. A bright feature appears detached 1000–1500 km above the limb on July 16.846 UT in the 953-nm filter, but it is not present in an image three minutes earlier. A possible interpretation is that the feature is visible by reflected sunlight and that the apparent detachment is due to the shadow of Jupiter on the plume.

Numerous reports, most of them received via the SL9 message center, indicate no detection of the impact of fragment B = 20. However, at the W. M. Keck Observatory, Mauna Kea, I. de Pater, J. Graham and G. Jernigan clearly recorded a faint plume in L (3.27–3.44 μm) at the expected position, starting on July 17.122 UT, fading around 17.134; and monitoring by S. Miller, M.-F. Jagod, T. Geballe and T. Brooke of the H_{3}^{+} ionospheric lines at 3.5 μm with the CGS4 échelle spectrometer on UKIRT showed a fivefold brightening of the emission around July 17.118, fading until 17.181.

The SL9 message center also relays that an impact associated with fragment C = 19 was detected by P. McGregor and M. Allen at 2.34 μm using the CASPIR infrared camera on the Australian National University 2.3-m telescope and by D. Crisp, V. Meadows, S. Lumsden and S. Lee using IRIS on the 3.9-m Anglo-Australian Telescope at Siding Spring. The event, seen at the limb on July 17.306 UT, brightened appreciably during the first five minutes and was still visible after one hour. A weaker event on July 17.267 has been identified by M. Brown as a probable recurrence of impact A after one Jupiter rotation.

The NASA/IRTF Comet Collision Science Team reports that NSFCAM 2.248-μm observations beginning on July 17.285 UT did not record any obvious flash on Io and Europa, but the data are yet to be photometrically reduced. On July 17.302 both the remnant of the A impact (with a surface brightness similar to that of the south polar hood) and a dim spot from fragment C were detected. By July 17.304 site C was considerably brighter than site A, but by July 17.311 site C had faded to about site A’s brightness, and it continued to fade until about July 17.319.

1994 July 17 Brian G. Marsden